OpenStudy (anonymous):

Solve. 6x2 + 8x = 14

OpenStudy (anonymous):

6x^2 + 8x = 14 => 3x^2 + 4x = 7 => 3x^2 + 4x - 7 = 0 Solve it using quadratic formula

OpenStudy (anonymous):

You can also solve it by factorization. \(3x^2+4x-7=(3x+7)(x-1)=0 \implies x=\frac{-7}{3}\) or \(x= 1\).

OpenStudy (anonymous):

Possible intermediate steps: 6 x^2+8 x = 14 Divide both sides by 6: x^2+(4 x)/3 = 7/3 Add 4/9 to both sides: x^2+(4 x)/3+4/9 = 25/9 Factor the left hand side: (x+2/3)^2 = 25/9 Take the square root of both sides: abs(x+2/3) = 5/3 Eliminate the absolute value: x+2/3 = -5/3 or x+2/3 = 5/3 Subtract 2/3 from both sides: x = -7/3 or x+2/3 = 5/3 Subtract 2/3 from both sides: x = -7/3 or x = 1